OpenStudy (anonymous):

solve each by completing the square. x^2+2x-3=0

OpenStudy (dangerousjesse):

Solve for x over the real numbers: \[x^2+2 x-3 = 0 \]Solve the quadratic equation by completing the square. Add 3 to both sides: \[x^2+2 x = 3 \]Take one half of the coefficient of x and square it, then add it to both sides. Add 1 to both sides: \[x^2+2 x+1 = 4 \]Factor the left hand side. Write the left hand side as a square: \[(x+1)^2 = 4 \]Eliminate the exponent on the left hand side. Take the square root of both sides: \[x+1 = 2~~ or ~~ x+1 = -2\]Look at the first equation: Solve for x. Subtract 1 from both sides: \[x = 1~~ or ~~ x+1 = -2 \]Look at the second equation: Solve for x. Subtract 1 from both sides Which leaves you with..?

OpenStudy (solomonzelman):

Didn't I delete it ? My thinking process when wrong. I DO recall that x²+2x+1 is a perfect square trinomial, and I know that 3 is 2 units away from 1, to get to the perfect square trinomial, but I don't get 1, I get -1..... this is the thinking error that I encountered. not like anyone is interested anyway, though.

OpenStudy (anonymous):

ok im getting confused a bit

OpenStudy (solomonzelman):

Oh, don't be afraid to hurt my feelings, I don't think I have them. Well... at least it's not like that I seem to have them :)

OpenStudy (anonymous):